### v3

@@
Sample Complexity for Constrained Markov Decision Process
@@
$\text{Jiashuo Jiang}^{{\dagger}}$, $\leavevmode\nobreak\ \text{Yinyu Ye}^{{\ddagger}\S}$
@@ 1 Introduction
Reinforcement learning (RL) is pivotal in the realm of dynamic decision-making under uncertainty, where the objective is to maximize total reward through ongoing interaction with and learning from an enigmatic environment. Markov Decision Processes (MDPs) are a prevalent framework for encapsulating environmental dynamics. MDPs have been instrumental in various domains, such as inventory management (Alvo et al. 2023), video gaming (Mnih et al. 2013), robotics (Kober et al. 2013), recommender systems (Shani et al. 2005), and beyond. Yet, they fall short in accommodating additional constraints that may influence the formulation of the optimal policy and the decision-maker’s engagement with the uncertain environment. Often, in MDP applications, there are stringent constraints on utilities or costs, emanating from areas like safe autonomous driving (Fisac et al. 2018), robotics (Ono et al. 2015), and financial management (Tamar et al. 2012). These constraints might also symbolize limitations on resources within resource allocation contexts with applications to revenue management (Jia et al. 2023, Jiang 2023, Li et al. 2023), inventory management and queueing systems (Chen et al. 2021). Constrained MDPs (CMDPs), as introduced in Altman (1999), enhance MDPs to factor in constraints affecting long-term policy results. In CMDPs, the decision-maker aims to optimize cumulative rewards while adhering to these constraints. Our paper focuses on CMDPs, and we aim to develop efficient algorithmic solutions.
The significance of RL in CMDP contexts has garnered substantial attention in recent years. A variety of methods for tackling CMDPs have been developed, including the primal-dual technique (e.g. Efroni et al. (2020)) which leverages the Lagrangian dual of CMDPs and implements an online learning strategy for the iterative update of dual variables. Other methods encompass constrained optimization (Achiam et al. 2017, Sun et al. 2024), the Lyapunov technique (Chow et al. 2018), among others. Previous research has established minimax bounds for CMDPs, representing the optimal regret that can be achieved for the most challenging problem within a specific class of problems. Nonetheless, these minimax regret bounds and worst-case scenarios can be overly conservative, leading to a disconnect between theoretical guarantees and practical performance for specific problem instances. A more tailored approach is warranted—one that ensures great performance on every single problem instance and furnishes problem-dependent guarantees. Our research takes the first step towards deriving optimal problem-dependent guarantees for CMDP problems.
@@ 1.1 Main Results and Contributions
In the exploration of the $\gamma$-discounted infinite horizon CMDP problem within a tabular context, we leverage a generative model that (as studied in Kearns et al. (2002), Kakade (2003), Li et al. (2024)), for each state-action pair, yields a sample from the unknown stochastic reward, costs, and state transitions. The main result of our research is the introduction of a novel algorithm that promises a $O(\kappa\cdot\log N)$ regret upper bound for both the cumulative reward and constraint violations, where $N$ is the number of queries to the generative model and $\kappa$ is a problem-specific parameter, yet independent of $N$. This translates into a $\tilde{O}\left(1/\varepsilon\right)$ sample complexity bound, where $\tilde{O}(\cdot)$ hides a term polynomial in $\log(1/\varepsilon)$. Note that the state-of-the-art sample complexity bound under the worst-case scenarios is $O(1/\varepsilon^{2})$, which has been established in a series of work (Wang 2020, Jin and Sidford 2020, Efroni et al. 2020). Though the $O(1/\varepsilon)$ or better iteration complexity has been achieved in Li et al. (2021b), Liu et al. (2021b), Zhou et al. (2022), Gladin et al. (2023), it comes with a sample complexity bound no better than $O(1/\varepsilon^{2})$. Our algorithm enjoys a sample complexity bound that has a better dependency in terms of $\varepsilon$. To achieve these problem-dependent guarantees, we develop several elements.
@@
First, we develop new characterizations of the problem instance hardness for CMDP problems. Note that a key component for achieving instance-dependent bounds is to characterize the “hardness” of the underlying problem instance. That is, we need to identify a positive gap to separate the optimal policy from the sub-optimal policies for a particular problem instance. The importance of identifying such a gap has been demonstrated in instance-optimal learning for multi-arm-bandit problems (e.g. Lai and Robbins (1985)) and reinforcement learning problems (e.g. Auer et al. (2008)), among others. The gap characterizes the hardness of the underlying problem instance (e.g. Simchowitz and Jamieson (2019)). For CMDP, identifying such a gap is non-trivial because the optimal policies for CMDP are randomized policies (Altman 1999). Then, the policies can be represented by distributions over the action set and the sub-optimal policies can be arbitrarily close to the optimal policy as long as the corresponding sub-optimal distributions converge to the optimal one. To tackle this problem, we show that the feasible region for the policies can be represented as a polytope and we only need to focus on the corner points of this polytope to find an optimal policy. Therefore, the hardness can simply be characterized as the distance between the optimal corner point and the sub-optimal corner point, as illustrated in detail in Section 3.1. This is the first characterization of problem instance hardness for CMDP problems.
@@
Then, we devise a new algorithmic framework to analyze CMDP problems, inspired by the online packing/linear programming (LP) literature (Agrawal et al. 2014, Kesselheim et al. 2014, Li and Ye 2022, Li et al. 2021c). Specifically, we utilize a linear programming reformulation of the CMDP problem, where policies are delineated via occupancy measures (Altman 1999). The optimal policy emerges from the LP’s solution; however, the indeterminate model parameters mean the LP cannot be solved directly but must be approached online as we obtain more samples from the generative model. Each generative model query leads to solving an empirical LP with accrued samples, and our final policy is derived from averaging these solutions—an approach akin to the methodology in online LP. A critical feature of our algorithm is the adaptiveness of the LP constraints’ right-hand side to the input samples, a technique demonstrated to achieve logarithmic regret in online LP literature, which we now apply to CMDP problems.
@@
Furthermore, we extend our contributions to the online LP literature. The prevalent assumption of non-degeneracy in online LP, which postulates a unique LP solution, is not a requirement in our approach. We introduce an innovative method to identify and consistently apply one optimal LP basis, even when multiple bases exist. Aligning with a single optimal basis—as opposed to requiring uniqueness—affords the same benefits and leads to logarithmic regret. To delve into details, our algorithm enhances performance through two new components. The first is a systematic procedure to pinpoint a particular optimal LP basis from potentially several (outlined in Algorithm 1), diverging from the traditional literature that presupposes a sole optimal solution and basis. The second component is a resolving procedure that capitalizes on the structure of the identified optimal basis. Once an optimal basis is selected, we can determine the corresponding optimal solution by solving a set of linear equations, which involves only the non-zero basic variables and the active constraints. This resolving process transforms the algorithmic idea of resolving an LP, as seen in previous works (Agrawal et al. 2014, Bumpensanti and Wang 2020, Li and Ye 2022, Vera and Banerjee 2021, Li et al. 2021c), into solving specific linear equations based on the chosen optimal basis (detailed in Algorithm 2). Previous research relies on the non-degeneracy assumption to stabilize the optimal basis throughout the resolving process. However, our two innovative elements enable our algorithm to consistently adhere to one optimal basis, thus removing the non-degeneracy assumption. This not only simplifies the algorithm’s execution (from resolving LP to resolving linear equations) but also extends its applicability, freeing it from assumptions that have traditionally limited the scope of online LP techniques.
@@
We finally discuss the dependency of our sample complexity bound on problem parameters other than $\varepsilon$. We restrict to the MDP context without resource constraints. Denote by $\mathcal{S}$ the state set and $\mathcal{A}$ the action set. We show a sample complexity bound of $O\left(\frac{|\mathcal{S}|^{4}\cdot|\mathcal{A}|}{(1-\gamma)^{4}\cdot\Delta}\cdot\frac{\log^{2}(1/\varepsilon)}{\varepsilon}\right)$, where $\Delta$ is the constant that represents the hardness of the underlying problem instance. Compared to the optimal worst-case sample complexity $O\left(\frac{|\mathcal{S}|\cdot|\mathcal{A}|}{(1-\gamma)^{3}\cdot\varepsilon^{2}}\right)$ that is achieved in a series of work (Sidford et al. 2018, Wainwright 2019, Wang 2020, Agarwal et al. 2020, He et al. 2021), our bound has a worse dependency over $|\mathcal{S}|$ and $1-\gamma$. This is due to our algorithm being LP-based and the dimension of the LP ($|\mathcal{S}|$ and $1-\gamma$) will influence our final bounds. However, our bound enjoys a better dependency in terms of $\varepsilon$. For the general CMDP problem, our bound will depend additionally on the conditional number of the constraint matrix in the LP formulation, which is a byproduct of the resolving LP heuristics (Vera and Banerjee 2021, Li et al. 2021c). However, our $\tilde{O}(1/\varepsilon)$ sample complexity bound depends polynomially on other parameters including $|\mathcal{S}|$, $|\mathcal{A}|$, $1-\gamma$, and the number of constraints. Partial results of this paper have been accepted at NeurIPS 2024 conference. Compared with the conference version, the current paper not only enriches the interpretations significantly and includes complete detailed proofs, but also includes the following key extensions.
@@
Extension to the finite horizon episodic setting. We extend our approaches and results to the finite horizon setting, where there are multiple periods in a horizon and the horizon can repeated by multiple times to learn the optimal policy. Note that the optimal policy under the finite horizon setting can still be written as a LP. We directly extend the online LP approach to handle the finite horizon setting and obtain similar results. To be specific, we show that in order to obtain an $\epsilon$-approximate policy where we approximate the optimal reward by a $\varepsilon$ gap and allows at most a $\varepsilon$ constraint violation, we need at most $\tilde{O}(1/\varepsilon)$ number of samples. The details are given in Section 7.1.
@@
Extension to the online and offline learning settings. We consider more involved sampling procedure beyond assuming generative model. Note that in practice, the agent usually interacts with the true environment (online learning setting), where the agent can only obtain the information regarding the state it is currently visiting and the action it is currently taking, or interacts with an offline simulator (offline learning setting), where there is a fixed historical data set which may be collected using some other algorithms. Different from the generative model where we can obtain information for arbitrary states and actions, the online and offline learning settings put further restrictions on the freedom of obtaining samples and require additional considerations when designing algorithms. However, we show that our approach can still be directly extended to handle the online and offline learning settings, and we are also able to derive the similar $\tilde{O}(1/\varepsilon)$ sample complexity. The details are presented in Section 7.2.
@@
In summary, our work presents a new algorithm for CMDP problems to achieve the problem-dependent logarithmic regret or $\tilde{O}(1/\varepsilon)$ sample complexity bound, for the first time in the literature. Firstly, we have pioneered the application of the online LP framework to analyze CMDP problems. Our algorithm operates in the primal space, in contrast to the primal-dual framework associated with worst-case guarantees (e.g. Ding et al. (2021)). Secondly, our novel approach to online LP results in logarithmic regret without the non-degeneracy assumption, expanding its applicability and providing new insights into both CMDP and online LP problems. We finally explore extensions to the finite horizon setting, as well as the online and offline learning settings, and obtain similar $\tilde{O}(1/\varepsilon)$ sample complexity bounds. These extensions demonstrate the generality of our approach.
@@ 1.2 Further Related Work
Deriving the (logarithmic) problem-dependent guarantees is a long-standing topic in online decision-making (Lai and Robbins 1985, Graves and Lai 1997). Notably, there has been a recent trend of deriving (logarithmic) problem-dependent guarantees for reinforcement learning problems, where a series of work (Zanette and Brunskill 2019, Simchowitz and Jamieson 2019, Mou et al. 2020, Foster et al. 2020, Dann et al. 2021, Bhandari and Russo 2021, Wagenmaker et al. 2022, Wagenmaker and Foster 2023, Xu and Zeevi 2024, Duan and Wainwright 2024, Chen et al. 2024) has proposed various ways to characterize the hardness of the problem instance. A prevalent way to characterize the hardness is to define a constant that represents the policy sub-optimality gap between the optimal policy and a policy that takes a sub-optimal action given a state. Then, the logarithmic guarantees depend on this constant. Compared to the previous approaches and guarantees, our guarantees adopt this idea of policy sub-optimal gap. Moreover, our approach can handle resource or safety constraints that need to be satisfied by the policy. In this way, our approach generalizes to multi-objective (safe) RL problems. Safe reinforcement learning problems have been investigated extensively under various settings in a series of work (Liu et al. 2021a, Wei et al. 2021, Ding et al. 2021, Li et al. 2021b, Ghosh et al. 2022, Chen et al. 2022, Ding and Lavaei 2023, Germano et al. 2023, Gladin et al. 2023), where the $O(\sqrt{T})$ regret bounds or $O(1/\varepsilon^{2})$ sample complexity bounds are proved.
Our method is motivated by the online LP literature. The online LP problem (Molinaro and Ravi 2014, Agrawal et al. 2014, Gupta and Molinaro 2014) covers a wide range of applications through different ways of specifying the underlying LP, including secretary problem (Ferguson 1989), online knapsack problem (Arlotto and Xie 2020), quantity-based network revenue management problem (Jasin and Kumar 2012), network routing problem (Buchbinder and Naor 2009), matching problem (Mehta et al. 2007), etc. Notably, the problem has been studied under either (i) the stochastic input model where the coefficient in the objective function, together with the corresponding column in the constraint matrix is drawn from an unknown distribution $\mathcal{P}$, or (ii) the random permutation model where they arrive in a random order (Molinaro and Ravi 2014, Agrawal et al. 2014, Kesselheim et al. 2014, Gupta and Molinaro 2014). With an additional non-degeneracy assumption, the logarithmic regret has been derived for the quantity-based network revenue management problem (Jasin and Kumar 2012, Jasin 2014), the online LP problem (Li and Ye 2022), and the more general convex allocation problem (Ma et al. 2022). The non-degeneracy assumption has been relaxed in a series of work (Bumpensanti and Wang 2020, Vera and Banerjee 2021, Jiang et al. 2022, Wei et al. 2023, Ao et al. 2024, Banerjee and Freund 2024) under various settings.

### v4

@@
Approaching Constrained Markov Decision Process
@@
$\text{Jiashuo Jiang}^{{\dagger}}$, $\penalty 10000\ \text{Yinyu Ye}^{{\ddagger}\S}$
@@ 1 Introduction
Reinforcement learning (RL) is pivotal in the realm of dynamic decision-making under uncertainty, where the objective is to maximize total reward through ongoing interaction with and learning from an enigmatic environment. Markov Decision Processes (MDPs) are a prevalent framework for encapsulating environmental dynamics. MDPs have been instrumental in various domains, such as inventory management (alvo2023neural), video gaming (mnih2013playing), robotics (kober2013reinforcement), recommender systems (shani2005mdp), and beyond. Yet, they fall short in accommodating additional constraints that may influence the formulation of the optimal policy and the decision-maker’s engagement with the uncertain environment. Often, in MDP applications, there are stringent constraints on utilities or costs, emanating from areas like safe autonomous driving (fisac2018general), robotics (ono2015chance), and financial management (tamar2012policy). These constraints might also symbolize limitations on resources within resource allocation contexts with applications to revenue management (jia2023online, jiang2023constant, li2023revenue), inventory management and queueing systems (chen2021primal). Constrained MDPs (CMDPs), as introduced in altman1999constrained, enhance MDPs to factor in constraints affecting long-term policy results. In CMDPs, the decision-maker aims to optimize cumulative rewards while adhering to these constraints. Our paper focuses on CMDPs, and we aim to develop efficient algorithmic solutions.
The significance of RL in CMDP contexts has garnered substantial attention in recent years. A variety of methods for tackling CMDPs have been developed, including the primal-dual technique (e.g. efroni2020exploration) which leverages the Lagrangian dual of CMDPs and implements an online learning strategy for the iterative update of dual variables. Other methods encompass constrained optimization (achiam2017constrained, sun2024constrained), the Lyapunov technique (chow2018lyapunov), among others. Previous research has established minimax bounds for CMDPs, representing the optimal regret that can be achieved for the most challenging problem within a specific class of problems. Nonetheless, these minimax regret bounds and worst-case scenarios can be overly conservative, leading to a disconnect between theoretical guarantees and practical performance for specific problem instances. A more tailored approach is warranted—one that ensures great performance on every single problem instance and furnishes problem-dependent guarantees. Our research takes the first step towards deriving optimal problem-dependent guarantees for CMDP problems.
@@ 1.1 Main Results and Contributions
[…]
@@
First, we develop new characterizations of the problem instance hardness for CMDP problems. Note that a key component for achieving instance-dependent bounds is to characterize the “hardness” of the underlying problem instance. That is, we need to identify a positive gap to separate the optimal policy from the sub-optimal policies for a particular problem instance. The importance of identifying such a gap has been demonstrated in instance-optimal learning for multi-arm-bandit problems (e.g. lai1985asymptotically) and reinforcement learning problems (e.g. auer2008near), among others. The gap characterizes the hardness of the underlying problem instance (e.g. simchowitz2019non). For CMDP, identifying such a gap is non-trivial because the optimal policies for CMDP are randomized policies (altman1999constrained). Then, the policies can be represented by distributions over the action set and the sub-optimal policies can be arbitrarily close to the optimal policy as long as the corresponding sub-optimal distributions converge to the optimal one. To tackle this problem, we show that the feasible region for the policies can be represented as a polytope and we only need to focus on the corner points of this polytope to find an optimal policy. Therefore, the hardness can simply be characterized as the distance between the optimal corner point and the sub-optimal corner point, as illustrated in detail in Section 3.1. This is the first characterization of problem instance hardness for CMDP problems.
@@
Then, we devise a new algorithmic framework to analyze CMDP problems, inspired by the online packing/linear programming (LP) literature (agrawal2014dynamic, kesselheim2014primal, li2022online, li2021symmetry). Specifically, we utilize a linear programming reformulation of the CMDP problem, where policies are delineated via occupancy measures (altman1999constrained). The optimal policy emerges from the LP’s solution; however, the indeterminate model parameters mean the LP cannot be solved directly but must be approached online as we obtain more samples from the generative model. Each generative model query leads to solving an empirical LP with accrued samples, and our final policy is derived from averaging these solutions—an approach akin to the methodology in online LP. A critical feature of our algorithm is the adaptiveness of the LP constraints’ right-hand side to the input samples, a technique demonstrated to achieve logarithmic regret in online LP literature, which we now apply to CMDP problems.
@@
Furthermore, we extend our contributions to the online LP literature. The prevalent assumption of non-degeneracy in online LP, which postulates a unique LP solution, is not a requirement in our approach. We introduce an innovative method to identify and consistently apply one optimal LP basis, even when multiple bases exist. Aligning with a single optimal basis—as opposed to requiring uniqueness—affords the same benefits and leads to logarithmic regret. To delve into details, our algorithm enhances performance through two new components. The first is a systematic procedure to pinpoint a particular optimal LP basis from potentially several (outlined in Algorithm 2), diverging from the traditional literature that presupposes a sole optimal solution and basis. The second component is a resolving procedure that capitalizes on the structure of the identified optimal basis. Once an optimal basis is selected, we can determine the corresponding optimal solution by solving a set of linear equations, which involves only the non-zero basic variables and the active constraints. This resolving process transforms the algorithmic idea of resolving an LP, as seen in previous works (agrawal2014dynamic, bumpensanti2020re, li2022online, vera2021bayesian, li2021symmetry), into solving specific linear equations based on the chosen optimal basis (detailed in Algorithm 3). Previous research relies on the non-degeneracy assumption to stabilize the optimal basis throughout the resolving process. However, our two innovative elements enable our algorithm to consistently adhere to one optimal basis, thus removing the non-degeneracy assumption. This not only simplifies the algorithm’s execution (from resolving LP to resolving linear equations) but also extends its applicability, freeing it from assumptions that have traditionally limited the scope of online LP techniques.
@@
We finally discuss the dependency of our sample complexity bound on problem parameters other than $\varepsilon$. We restrict to the MDP context without resource constraints. Denote by $\mathcal{S}$ the state set and $\mathcal{A}$ the action set. We show a sample complexity bound of $O\left(\frac{|\mathcal{S}||\mathcal{A}|}{(1-\gamma)^{4}\cdot\Delta}\cdot\frac{\log^{2}(1/\varepsilon)}{\varepsilon}\right)$, where $\Delta$ is the constant that represents the hardness of the underlying problem instance. Compared to the optimal worst-case sample complexity $O\left(\frac{|\mathcal{S}|\cdot|\mathcal{A}|}{(1-\gamma)^{3}\cdot\varepsilon^{2}}\right)$ that is achieved in a series of work (sidford2018near, wainwright2019variance, wang2020randomized, agarwal2020model, he2021nearly), our bound has a worse dependency over $1-\gamma$. This is due to our algorithm being LP-based and the dimension of the LP (can be bounded by $1-\gamma$) will influence our final bounds. However, our bound enjoys a better dependency in terms of $\varepsilon$. For the general CMDP problem, our bound will depend additionally on the conditional number of the constraint matrix in the LP formulation, which is a byproduct of the resolving LP heuristics (vera2021bayesian, li2021symmetry). However, our $\tilde{O}(1/\varepsilon)$ sample complexity bound depends polynomially on other parameters including $|\mathcal{S}|$, $|\mathcal{A}|$, $1-\gamma$, and the number of constraints.
@@
Extension to the finite horizon episodic setting. We extend our approaches and results to the finite horizon setting, where there are multiple periods in a horizon and the horizon can repeated by multiple times to learn the optimal policy. Note that the optimal policy under the finite horizon setting can still be written as a LP. We directly extend the online LP approach to handle the finite horizon setting and obtain similar results. To be specific, we show that in order to obtain an $\varepsilon$-approximate policy where we approximate the optimal reward by a $\epsilon$ gap and allows at most a $\varepsilon$ constraint violation, we need at most $\tilde{O}(1/\varepsilon)$ number of samples. The details are given in Section 7.1.
@@
Extension to the off-policy learning settings. We consider more involved sampling procedure beyond assuming generative model. Note that in practice, the agent usually interacts with an offline simulator (offline learning setting), where there is a fixed historical data set which may be collected using some other algorithms. Different from the generative model where we can obtain information for arbitrary states and actions, the offline learning settings put further restrictions on the freedom of obtaining samples and require additional considerations when designing algorithms. However, we show that our approach can still be directly extended to handle the offline learning settings, and we are also able to derive the similar $\tilde{O}(1/\varepsilon)$ sample complexity. The details are presented in Section 7.2.
Extension to the on-policy learning settings. We consider a more involved sampling procedure when the agent interacts with the true environment (online learning setting), where the agent can only obtain the information regarding the state it is currently visiting and the action it is currently taking. Different from the off-policy learning setting where we can use an additional behavior policy to collect samples, which could be different from the target policy, in the on-policy learning setting, we need to collect samples using the output generated by our algorithm. Therefore, our algorithm not only needs to approximate the optimal policy, but also needs to enjoys the property to do exploration automatically by visiting the desired states as much as possible. We further modify our algorithm and show that our approach can still handle the on-policy learning setting and achieve a similar $\tilde{O}(1/\varepsilon)$ sample complexity. The details are presented in Section 7.3.
@@
In summary, our work presents a new algorithm for CMDP problems to achieve the problem-dependent logarithmic regret or $\tilde{O}(1/\varepsilon)$ sample complexity bound, for the first time in the literature. Firstly, we have pioneered the application of the online LP framework to analyze CMDP problems. Our algorithm operates in the primal space, in contrast to the primal-dual framework associated with worst-case guarantees (e.g. ding2021provably). Secondly, our novel approach to online LP results in logarithmic regret without the non-degeneracy assumption, expanding its applicability and providing new insights into both CMDP and online LP problems. We finally explore extensions to the finite horizon setting, as well as the online and offline learning settings, and obtain similar instance-dependent guarantees. These extensions demonstrate the generality of our approach.
@@ 1.2 Further Related Work
Deriving the (logarithmic) problem-dependent guarantees is a long-standing topic in online decision-making (lai1985asymptotically, graves1997asymptotically). Notably, there has been a recent trend of deriving (logarithmic) problem-dependent guarantees for reinforcement learning problems, where a series of work (zanette2019tighter, simchowitz2019non, foster2020instance, dann2021beyond, bhandari2021linear, wagenmaker2022beyond, wagenmaker2023instance, xu2024towards, duan2024taming, chen2024landscape) has proposed various ways to characterize the hardness of the problem instance. A prevalent way to characterize the hardness is to define a constant that represents the policy sub-optimality gap between the optimal policy and a policy that takes a sub-optimal action given a state. Then, the logarithmic guarantees depend on this constant. Compared to the previous approaches and guarantees, our guarantees adopt this idea of policy sub-optimal gap. Moreover, our approach can handle resource or safety constraints that need to be satisfied by the policy. In this way, our approach generalizes to multi-objective (safe) RL problems. Safe reinforcement learning problems have been investigated extensively under various settings in a series of work (e.g. liu2021learning), where the $O(\sqrt{T})$ regret bounds or $O(1/\varepsilon^{2})$ sample complexity bounds are proved.
Our method is motivated by the online LP/resource allocation literature. The online LP problem (molinaro2014geometry, agrawal2014dynamic, gupta2014experts) covers a wide range of applications through different ways of specifying the underlying LP, including secretary problem (ferguson1989solved), online knapsack problem (arlotto2020logarithmic), quantity-based network revenue management problem (jasin2012re), network routing problem (buchbinder2009online), matching problem (mehta2007adwords), etc. Notably, the problem has been studied under either (i) the stochastic input model where the coefficient in the objective function, together with the corresponding column in the constraint matrix is drawn from an unknown distribution $\mathcal{P}$, or (ii) the random permutation model where they arrive in a random order (molinaro2014geometry, agrawal2014dynamic, kesselheim2014primal, gupta2014experts). With an additional non-degeneracy assumption, the logarithmic regret has been derived for the quantity-based network revenue management problem (jasin2012re, jasin2014reoptimization), the online LP problem (li2022online), and the more general convex allocation problem (ma2022optimal). The non-degeneracy assumption has been relaxed in a series of work (bumpensanti2020re, vera2021bayesian, jiang2022degeneracy, wei2023constant, ao2024online, banerjee2024good) under various settings.